### v2

# Attributing the point symmetric structure of core-collapse supernova remnant N132D to the jittering jets explosion mechanism

Abstract
I identified a point-symmetric morphology in the core-collapse supernova (CCSN) remnant (CCSNR) N132D, composed of two symmetry axes: the short symmetry axis extending from the northwest ear and through the center of the iron-rich emission on the other side, and the second along the long dimension of N132D and coincides with the extension of the central oxygen-rich gas to the northeast. Namely, the point-symmetry of the outer zones of CCSNR N132D correlates with that of the oxygen-rich gas near the center. The surrounding gas cannot shape the inner oxygen-rich material, implying that the point-symmetric morphology is a property of the explosion mechanism, as predicted by the jittering jets explosion mechanism (JJEM). The oxygen-rich material is known to be in a torus. According to the JJEM, an energetic pair of opposite jets, more or less perpendicular to the plane of the torus, has shaped the torus; this pair is along the short symmetry axis. Another energetic pair, perpendicular to the first one, shaped the elongated, large-scale structure of CCSNR N132D. I discuss how the JJEM accounts for two perpendicular pairs of jets and the unequal jets in each pair. CCSNR N132D is the fifteenth CCSNR with an identified point-symmetric morphology attributed to the JJEM. Because the neutrino-driven mechanism cannot explain such morphologies, this study further strengthens the claim that the JJEM is the primary explosion mechanism of CCSNe.
keywords:
supernovae: general – stars: jets – ISM: supernova remnants – stars: massive1 Introduction
Studies in the last two years have been discussing two competing core-collapse supernova (CCSN) explosion mechanisms, the delayed neutrino explosion mechanism (e.g., Andresenetal2024; BoccioliFragione2024; Burrowsetal2024kick; JankaKresse2024; vanBaaletal2024; WangBurrows2024; Bambaetal2025CasA; Bocciolietal2025; EggenbergerAndersenetal2025; Huangetal2025; Imashevaetal2025; Laplaceetal2025; Maltsevetal2025; Maunderetal2025; Morietal2025; Mulleretal2025; Nakamuraetal2025; SykesMuller2025; Janka2025; ParadisoCoughlin2025; Vinketal2025; WangBurrows2025) and the jittering jets explosion mechanism (JJEM), which is the frame of the present study.
The JJEM asserts that the primary explosion mechanism of the majority, and likely all, CCSNe involves pairs of jets (e.g., Soker2010; PapishSoker2011). The magnetorotational explosion mechanism is much older, but significantly differs from the JJEM. In the magnetorotational explosion mechanism, one pair of jets along a fixed axis explodes the star. To maintain a fixed angular momentum axis along which the accretion disk launches a pair of jets, the progenitor of the CCSN must be rapidly rotating (e.g., Shibagakietal2024; ZhaMullerPowell2024; Shibataetal2025 and references to much older papers therein). The requirement for a rapidly rotating pre-collapse core makes this mechanism rare. Therefore, the magnetorotational explosion mechanism attributes most CCSNe to the neutrino-driven mechanism, and I classify it as part of the neutrino-driven mechanism.
According to the JJEM, pairs of jittering jets that intermittent accretion disks (or belts) around the newly born neutron star (NS) launch on a timescales of explode all CCSNe (for a list of the parameters of the JJEM see Soker2025Learning); time is measured from shock bounce, i.e., the formation time of the shock at the newly born NS. In rare cases when the pre-explosion core is rapidly rotating, the jittering of the pairs of jets is at very small angles around the pre-explosion angular momentum axis. Practically, it is a fixed-axis explosion, some of which end with a black hole remnant. The variations in the axes of the pairs of jittering jets might be fully or partially stochastic. The two jets in a pair might substantially differ in their opening angle and power (e.g., Bearetal2025Puppis), and might be at an angle smaller than , namely, not exactly opposite (e.g., Shishkinetal2025S147).
Angular momentum fluctuations in the convection zones of the collapsing core that instabilities above the NS (for instabilities turbulence in the gain region, see, e.g., Abdikamalovetal2016; KazeroniAbdikamalov2020; Buelletetal2023) amplify cause the stochastic variations in the angular momentum axis directions of the accreted gas that launches the jets along these axes (e.g., GilkisSoker2014; GilkisSoker2016; ShishkinSoker2021; ShishkinSoker2023; WangShishkinSoker2024). Neutrino heating plays a role by adding energy to the jets (Soker2022nu). However, most of the energy originates from the jittering jets, which serve as the primary driver of the explosion.
In a recent review (Soker2024UnivReview), I summarized the state of debate between these two alternative explosion mechanisms as of 2024. The main challenge of the JJEM is to demonstrate with three-dimensional numerical simulations the formation of intermittent accretion disks during the explosion process. Such simulations should be with very high resolution, include magnetic fields, and the correct pre-collapse core convection. The next step of launching the jets by these intermittent accretion disks seems to be beyond the present capabilities of numerical simulations Soker2025Learning. The neutrino-driven mechanism encounters problems and difficulties that the community often overlooks. I list four issues.
(1) There are qualitative and quantitative disagreements between simulations of different research groups (e.g., Janka2025), like on which stellar models explode and which do not.
(2) Some modelings of CCSNe deduce explosion energies of (e.g., Moriyaetal2025), more than the neutrino-driven mechanism can supply. Many magnetar models of superluminous CCSNe require explosion energies of that imply explosion by jets (e.g., SokerGilkis2017; Kumar2025), a conclusion overlooked by most magnetar modelers.
(3) The neutrino mechanism predicts that some massive stars fail to explode, in contradiction with studies that suggest that there is only a small or no population of failed CCSNe (e.g., ByrneFraser2022; StrotjohannOfekGalYam2024; BeasoretalLuminosity2025; Healyetal2025). Bocciolietal2025 suggested that the neutrino-driven mechanism can overcome this difficulty.
(4) The most severe and challenging problem the neutrino-driven mechanism encounters is explaining the point-symmetric morphologies of CCSN remnants (CCSNRs; Section 2). This problem likely rules out the neutrino-driven mechanism as the primary mechanism of the explosion. On the other hand, the JJEM predicts that many CCSNRs exhibit point-symmetric morphologies, as recent three-dimensional hydrodynamical simulations demonstrate (Braudoetal2025). For this reason, identifying point-symmetric CCSNRs is of high significance in determining the explosion mechanism. In Section 3, I claim that SNR N132D is another CCSNR with a point-symmetric morphology that supports the JJEM.
Because many researchers are unaware of these difficulties or purposely ignore them in papers and review talks, the community overrates the neutrino-driven mechanism, as I discuss in the Summary of this study (Section 4).
2 Identifying point-symmetry by eyes
The pairs of opposite jets that explode the star in the JJEM might shape pairs of structural features in the descendant CCSNR. I emphasize that in most CCSNe these jets are not relativistic (Guettaetal2020 suggest that most CCSNe have no relativistic jets, contrary to gamma ray bursts that do have, e.g., Izzoetal2019; AbdikamalovBeniamini2025). Only a small number of pairs leave imprints on the descendant CCSNR; the other explode the core, lose their symmetry, and leave no clear geometrical imprint. Because the symmetry axes of the pairs of jets change their directions from one jet-launching episode to the next, the CCSNR’s pairs of structural features do not share the same axis, resulting in a point-symmetric morphology. The opposite structural features might include dense clumps, dense elongated structures termed filaments, bubbles, which are faint structures closed and encircled by a brighter rim, lobes, which are bubbles with partial rims, and ears, which are protrusions from the main CCSNR shell with decreasing cross-sections away from the center.
Studies identified point-symmetric morphologies in 14 CCSNRs and discussed these in the frame of the JJEM. Some of these CCSNRs possess clear point-symmetrical morphological features, like N63A (Soker2024CounterJet), the Vela CCSNR (Soker2023SNRclass; SokerShishkin2025Vela), Cassiopeia A (BearSoker2025), and the Crab Nebula (ShishkinSoker2025Crab), and others with less secure point-symmetric morphologies, like CTB 1 (BearSoker2023RNAAS). Other CCSNRs and the studies that attributed their morphologies to the JJEM are SNR 0540-69.3 (Soker2022SNR0540), the Cygnus Loop (ShishkinKayeSoker2024), SN 1987A (Soker2024NA1987A; Soker2024Keyhole), G321.3–3.9 (Soker2024CF; ShishkinSoker2025G321), G107.7-5.1 (Soker2024CF), W44 (Soker2024W44), Puppis A (Bearetal2025Puppis), SNR G0.9+0.1 (Soker2025G0901), and S147 Shishkinetal2025S147. In some CCSNRs, the dust also presents point-symmetric morphologies (Soker2025Dust).
Several processes are likely to disrupt the point-symmetric morphology, such as instabilities and the NS natal kick velocity that occur during the explosion, as well as certain post-explosion processes. The latter include interaction with a circumstellar material (CSM) lost by the CCSN progenitor (e.g., Chiotellisetal2021; ChiotellisZapartasMeyer2024; Velazquezetal2023; Meyeretal2022; MeyerDetal2024), interaction with the interstellar medium (ISM; e.g., Wuetal2019; YanLuetal2020; LuYanetal2021; MeyerMelianietal2024), pulsar wind nebula if it exists, and heating processes such as reverse shock and radioactive decay. In many CCSNRs, the processes that smear the point-symmetric structures induced by the explosion make the identification of point-symmetric morphologies challenging or even impossible.
In this study of the SNR N132D and the earlier studies of point-symmetric CCSNRs, I have been utilizing my over thirty years of experience in classifying planetary nebulae by inspecting their jet-shaped morphologies. The visual-inspection classification of morphologies is a powerful common practice in classifying planetary nebulae (e.g., Balick1987; Chuetal1987; Sahaietal2007; Sahaietal2011), and AGN jets (e.g., Hortonetal2025), and has led to significant breakthroughs, particularly in establishing the major role of binary interaction in the shaping of planetary nebulae, and the role that jets play in many of these interacting binaries.
Although the visual-inspection method might seem completely subjective because it is not quantitative, it is not; it has a large objective component. Humans possess a remarkable ability to critically examine symmetry and identify departures from it by inspection alone, as facial symmetry is a key indicator of mate quality (e.g., Rhodes2006; Pinheiroetal2023 and reference therein;). In particular, our ability to identify fluctuating asymmetries, which are non-directional (random) deviations from perfect symmetry in bilaterally paired traits, is crucial as fluctuating asymmetries tend to reflect health problems (e.g., Rhodes2006). The majority of authors (and referees) made one of the most important decisions of their lives, if not the most important, i.e., choosing a mate, primarily through visual inspection, particularly by looking for symmetry and ruling out large fluctuating asymmetries. Recognition of symmetry is relatively automatic and consistent across cultures (e.g., Rhodes2006).
3 The point symmetric structure of SNR N132D
Numerous papers study the structure and morphology of SNR N132D (e.g., Lasker1978; Lasker1980; Hughes1987; Blairetal1994; DickelMilne1995AJ; Morseetal1995; Morseetal1996; Blairetal2000; Beharetal2001; Tappeetal2006; Tappeetal2012; XiaoAndChen2008; Bambaetal2018; Lawetal2020; Shardaetal2020; Suzukietal2020; Banovetzetal2023; Rhoetal2023; Xrism2024N132D; Fosteretal2025; LongX2025; Okadaetal2025). None of these studies examined the point-symmetric morphology and structure in the context of the JJEM, which is the goal of this study. Some studies have mentioned a possible jet as an explanation for the runaway knot in the southeast (e.g., VogtDopita2011), which I will refer to later.
Morseetal1995 first identified a point-symmetric structure of the fast O-rich ejecta of SNR N132D (Section 3.2); they did not term it a point-symmetric structure and did not connect it to jet shaping. Before turning to their results, I visually inspect some X-ray images to identify a global point-symmetric morphology (Section 3.1). In Section 3.3, I discuss the three-dimensional (3D) structure from VogtDopita2011 and Lawetal2020 in the frame of the point symmetric morphology and the JJEM.
3.1 The X-ray point-symmetric morphology
The recent study by Fosteretal2025 that maps iron emission, as I present in panel (a) of Figure 1, allows me to identify a symmetry axis, the short-symmetry axis of SNR N132D, extending from the northwest ear (NW ear) and through the center of the iron emission; this is the dashed-pale-blue line in panel (a). The X-ray image from the Chandra site, panel (b) of Figure 1, reveals a clear ear in the northwest. The protrusion on the opposite side is small and not distinct enough to be identified as a separate structural feature, although the two dents are. The new image by Fosteretal2025 presents a clear opposite structure to the northwest ear: a zone extending from the center to the southeast: the yellow-colored zones in panel (a) of Figure 1 that depict strong Fe K emission. The line along the long axis of the strong Fe K emission in the southeast crosses to the other side into the northwest ear. This is the short symmetry axis. The bright zones of other emission lines might be in different regions than the bright Fe K region, e.g., the S He, Fe He, and Fe Ly in the XRISM maps (Xrism2024N132D); these maps have lower spatial resolution than the XMM-Newton maps of Fosteretal2025. I use the bright zone of the Fe K line from Fosteretal2025 because this zone occupies a well-defined region that is bound by the two dents that I mark on panel (b) of Figure 1.
The iron emission Fe K is concentrated to the southeast. Xrism2024N132D find that the Fe Ly emission from SNR N132D is redshifted with a bulk velocity of . In panel (b) of Figure 1, I present an X-ray image from the Chandra site. This X-ray image shows three dents (bent inwards) on the outer boundary of the SNR, which I have marked with three yellow arrows. A fourth dent is inside the boundary (yellow-dashed arrow). I connect two pairs of dents with two solid red lines. The solid-orange double-sided arrow is the short-symmetry axis, i.e., at the same location as the dashed line on panel (a) of Figure 1. I added a double-sided arrow (dashed orange) for orientation: it is perpendicular to the short-symmetry axis and at the same length; the two double-sided arrows intersect at their centers. As immediately seen, SNR N132D is elongated in the northeast. The southwest end of the perpendicular double-sided arrow does not reach the boundary of the SNR, implying that the SNR is also elongated to the southwest.
Banovetzetal2023 measured the proper motion of the oxygen-rich ejecta of SNR N132D from two HST observations 16 years apart. They determined the center of expansion (CoE); in panel (c) of Figure 1, which I adapted from Banovetzetal2023, they mark this center with a yellow plus symbol. They also mark the two centers that Morseetal1995 identified, which I refer to in Section 3.2. I added to panel (c) the intersections of the two double-sided arrows and of the two red lines on panel (b) of Figure 1. The Chandra X-ray image of SNR N132D in panel (d) of Figure 1, adapted from Borkowskietal2007, shows the global structure of SNR N132D, including the centers of different structures from panel (c), the structures from panel (b), and the contours of the oxygen-rich material near the center (more on the oxygen-rich ejecta in Section 3.2 and 3.3).
While I based the short-symmetry axis, which is the double-sided arrows from southeast to the NW ear, on structural features, i.e., the NW ear and the iron-rich zone from Fosteretal2025, the dashed-orange double-sided arrow is simply a perpendicular line with the same length and through the center of the short-symmetry axis. A key feature for identifying point-symmetric morphology is that this perpendicular line aligns with the longest structure of the oxygen-rich ejecta near the center, indicating a clear association. However, I can also draw a line from a small ear on the southwest and through the center of the short-symmetry axis; this is the dotted-orange double-sided arrow. The southwest end of the dotted-orange double-sided arrow touches the edge of that ear, and its center is at the center of the short-symmetry axis. The dashed and dotted orange double-sided arrows are to each other.
I attribute the long extension of SNR N132D to a pair of opposite jets (or more). The angle of reflects the uncertainty in the location of the jets’ axis. Alternatively, the two jets were not exactly opposite to each other, but rather had an angle of between them. Such a ‘bent symmetry’, i.e., the two opposite sides are close to being at but not exactly, is observed in many planetary nebulae. Shishkinetal2025S147 suggest such bent opposite jets for CCSNR S147. I mark with solid arrows on panel (d) of Figure 1 the two jets I propose for this possibility.
The main result of this section is the identification of two axes that describe the large-scale structure, the short-symmetry axis and the perpendicular axis that extend along the long dimension of SNR N132D. I turn to show, Sections 3.2 and 3.3, that these axes correlate with those of the inner structure of the oxygen-rich ejecta.
3.2 Previously identified point-symmetric morphology
In Figure 2 I present two panels from Morseetal1995 of SNR N132D. Panel (a) presents in colors the velocity map by [O iii]5007 of fast oxygen-rich filaments and in gray the low-velocity oxygen emission. Panel (b) presents the O-rich filaments with the four lines that Morseetal1995 added to the image by connecting regions that show a symmetric distribution about a common center. They identified a point-symmetric morphological region in the center (although they do not use this term). They identified the intersection of these four lines as the center of the inner, fast, oxygen-rich filament distribution, which they marked by "" in panel (a). The "" symbol in panel (a) marks the center of the outer diffuse oxygen emission, which they determined by fitting an ellipse to that gas. These two centers also appear in panel (c) of Figure 1.
I added to panel (b) of Figure 2 the two double-sided arrows from Figure 1, that are the short-symmetry axis that I identify, and the perpendicular axis, which is along the long dimension of SNR N132D. From Figures 1 and 2 I notice the following. (1) Although the two axes I identify do not coincide with those of Morseetal1995, the perpendicular axis is also along the long dimension of the fast oxygen filaments in the center. (2) The center I identify by the two double-sided arrows (orange-dot on panels c and d of Figure 1) is close to the center of the inner oxygen filaments that Morseetal1995 identify.
I conclude that the outer ejecta and the inner ejecta share morphological features. This has significant implications for the shaping mechanism, as the CSM and ISM are unable to shape the inner ejecta. I discuss this and the perpendicular symmetry axes in Section 4.
3.3 The relation to the oxygen 3D structure
VogtDopita2011 map the [O III]5007 dynamics of SNR N132D and reconstruct its 3D structure. They found that the majority of the ejecta form a ring (torus-like) of in diameter. Lasker1980 already identified the ring structure of the oxygen-rich ejecta near the center. VogtDopita2011 speculated that the oxygen-rich ring is in the equatorial plane of a bipolar explosion and that the morphology is also strongly influenced by the CSM. In panel (a) of Figure 3, I present an image adapted from Rhoetal2023 that shows the Doppler shift measurements of Morseetal1995 on top of the general image of SNR N132D in the IR and X-ray. I added the three axes from panel (d) of Figure 1. The IR emission (diffuse blue regions) reveals material surrounding SNR N132D; however, I do not attribute the structure of the inner oxygen-rich ring (torus) to the CSM or ISM that are located outside the main SNR shell.
Additionally, VogtDopita2011 associated the fast, oxygen-rich runaway knot (RK; panel b of Figure 2) with a polar jet. I accept this interpretation that polar jets shaped the ring (torus) of oxygen-rich ejecta in the inner part of SNR N132D. I suggest that a pair of opposite jets along the short-symmetry axis participated in the explosion process of SNR N132D; there were several more pairs of jets according to the JJEM. In panel (b) of Figure 3, I present the three axes I identified in Figure 1 on top of an image adapted from Lawetal2020. The comparison in the panels of Figure 3 strengthens the relation I found from Figure 2 between the large-scale symmetry of SNR N132D and the symmetry of the oxygen-rich ejecta in the inner region.
The 3D structure reconstruction by Lawetal2020 is broadly consistent with the 3D geometry constructed by VogtDopita2011. Lawetal2020 find the majority of the bright oxygen ejecta to reside in a broken and distorted torus tilted to the plane of the sky and with a radius of , where is the distance to SNR N132D. The velocity along the line of sight is from to . They find the Doppler velocity of the runaway knot to be and its total space velocity , about twice the bulk velocity of the oxygen-rich ejecta of . Their finding supports the argument of VogtDopita2011 for a polar jet. Of high significance to my claim for the JJEM is Lawetal2020 finding that the runaway knot is nearly perpendicular to the torus plane and coincident with an X-ray emission spot (inset of panel (b) of Figure 3) that is substantially enhanced in silicon and somewhat in sulfur relative to the Large Magellanic Cloud and N132D’s bulk ejecta. The different composition from the rest of the ejecta shows that the runaway knot is not a result of an instability, e.g., a finger of Rayleigh-Taylor instability that keeps the same composition stratification of the rest of the ejecta; the runaway knot must be a separate ejection event, namely, a jet. This is the same argument on why the silicon-rich jet of Cassiopeia A cannot be an instability, and it is a jet launched during the explosion process (e.g., Soker2017IAUS).
Lawetal2020 find some departures from pure axi-symmetry of the torus in that the redshifted material is elevated above the midplane of the torus, while the blueshifted side is below the midplane. In the JJEM, this is part of the point-symmetric morphology that pairs of jets shape during the explosion process. Lawetal2020 also find that the runaway knot is to the normal to the torus.
Lawetal2020 identified a break in the torus; I present their identification in Figure 4. Examining some images from their analysis, I identify a counter break, i.e., a break opposite to the one they identified, as shown in panel (a) of Figure 4; it has a narrower opening. In the same panel, I draw a line from the break through the center and to the counter break that I identify here. As far as I can tell, this symmetry line through the breaks in the torus coincides on the plane of the sky with the perpendicular axis that I draw in Figures 1 and 2, up to the uncertainty in the exact location of the center of explosion. I suggest that a jet, one of two jets in a pair, shaped the break in the torus. The continuation of material outward on the sides of the break supports such an interpretation. An opposite jet shaped the counter break. On panel (b) of Figure 4, I added the perpendicular axis, but shifted to the center that Lawetal2020 uses. This axis is aligned along the extension of the oxygen-rich material to the northeast, highlighting the relationship between the large-scale symmetry and the oxygen-rich material.
Lawetal2020 compare the torus they reconstruct in SNR N132D with that of Cassiopeia A. I note that the torus of Cassiopeia A possesses a rich point-symmetric morphology (BearSoker2025). I here argue that the torus of SNR N132D also possesses a point-symmetric morphology. Lawetal2020 compared the runaway knot of SNR N132D to fast ejecta clumps in the Vela SNR, which is another CCSNR with a prominent and rich point-symmetric morphology that only the JJEM can account for (Soker2023SNRclass; SokerShishkin2025Vela).
The maps of N132D show some structures that are not part of a point-symmetric morphology, like the prominent bright filament extending south–southwest from the center in panels (b) and (d) of Figure 1. This is expected because several processes act to destroy point-symmetry: Instabilities at the explosion process as in the neutrino-driven simulations (see references in Section 1), one jet in a pair that is much more energetic than the opposite jet (Soker2024CounterJet; Bearetal2025Puppis), interaction with an asymmetrical CSM, the NS kick at explosion, nickel bubbles which are concentrations of nickel that when decay heat the gas and inflate bubbles, and binary star at explosion (see discussion by SokerShishkin2025Vela). This specific filament seems to result from unequal jets with instabilities that increased the asymmetry. This is for the future three-dimensional simulations of the JJEM.
This section solidifies my identification of a point-symmetric morphology in SNR N132D, and its relation to some properties in other point-symmetric CCSNRs.
4 Discussion and Summary
By visually inspecting images of CCSNR N132D, a method I justified in Section 2, I identified a point-symmetric morphology in CCSNR N132D, composed of two symmetry axes. The first one is the short symmetry axis, extending from the northwest ear and through the center of the iron-rich emission that Fosteretal2025 presented in a recent study (panel a of Figure 1). The second, the perpendicular axis on the plane of the sky, is an axis perpendicular to the short symmetry axis and passing through its center, which is also aligned along the long dimension of SNR N132D; it coincides with the extension of the oxygen-rich gas to the northeast (panel d of Figure 1, Figure 2 and 3, and panel b of Figure 4). The perpendicular axis seems to coincide with the line from the break in the torus that Lawetal2020 identify to the counter break that I identify (panel a of Figure 4). There might be a third axis: Lawetal2020 find the redshifted material in the torus to be somewhat above the midplane of the torus and the blueshifted side to be below the midplane. A line between these two regions, that is inclined to the plane of the torus, might define a third symmetry axis in SNR N132D. This deserves further study.
SNR N132D is surrounded by material (CSM and/or ISM), as indicated by the diffuse blue zones in panel (a) of Figure 3. SNR N132D interacts with the surrounding clouds (e.g., Dopitaetal2018; Sanoetal2020; HESSColl2021; Guetal2025N132D). This interaction affects the morphology of the outer regions, but cannot shape the inner oxygen-rich gas. The fact that the inner and outer morphologies correlate shows that the point symmetry is a result of the explosion.
I consider two energetic pairs of jets to shape the two symmetry axes; other, weaker pairs of jets that contributed to the explosion process are possible and likely. One pair of energetic jets is perpendicular to the plane of the oxygen-rich torus (Figure 4). This jet axis coincides with the short symmetry axis. The runaway knot (Figures 2 - 4) is a remnant of one of these two energetic jets. The unequal sides, an ear on the northwest and an iron-rich zone in the southeast (panel a of Figure 1), suggest that the jets were unequal in their power and/or opening angles. This deserves further study. The second pair of jets is perpendicular to the first pair of jets and is aligned along the long dimension of SNR N132D, coinciding with the long axis of the oxygen-rich material, as shown in all figures. The two jets were unequal, with the northeast jet being significantly more powerful, as indicated by the more extended SNR in that direction.
I considered the possibility that the two opposite jets along the long dimension of SNR N132D were not exactly opposite, i.e., a bent pair of jets. I mark the two suggested jets’ directions with solid-orange arrows in panel (d) of Figure 1; they are at to each other. Shishkinetal2025S147 suggested that one of the pairs of jets that exploded and shaped CCSNR S147 was a bent pair.
The JJEM accounts for the point-symmetric morphology of SNR N132D, as well as for the perpendicular two axes and the bent pair of jets. The perpendicular axes might result from the following two effects. In an early energetic pair of jets, one jet carries much more momentum (and energy) than the other. As a result of that, the NS acquires a kick velocity; this is the kick-BEAP (kick by early asymmetrical pair) mechanism (Bearetal2025Puppis; Shishkinetal2025S147). The kick velocity is along the axis of the jets. The kick velocity imparts an angular momentum component to the material that the NS accretes later from the collapsing core that is perpendicular to the kick velocity. If this component dominates, then the pair of jets that the accreted gas launches along its angular momentum axis tends to be perpendicular to the kick velocity (e.g., BearSoker2018kick), i.e., perpendicular to the axis of the first pair of jets. This effect of post-kick accretion might explain the avoidance of small angles between the NS kick velocity and the main jet axis in some CCSNRs (e.g., Soker2022SNR0540; BearSoker2023RNAAS). Many of the accretion episodes through the intermittent accretion disks in the JJEM last for a time scale that is shorter, or not much longer than the relaxation time of the accretion disk (Soker2024CounterJet). Because accretion of material with stochastic angular momentum and density fluctuations forms the accretion disk, the two sides of the unrelaxed accretion disk might be unequal in size and structure. Earlier studies of the JJEM followed (Soker2024CounterJet) and considered that opposite jets differ in their power and opening angle. I here add to the claim by Shishkinetal2025S147 that, in addition, the two jets might not be exactly opposite to each other, and the angle between them can be , e.g., in SNR S147.
One of the future steps will be to conduct heavy three-dimensional hydrodynamical simulations of the expansion of the jets through the inner core, i.e., in the silicon and oxygen layers, and follow the nucleosynthesis that the jets induce. These simulations will find the kinematics of the different elements. They will account for the dynamics, e.g., the finding that the velocity dispersions of the sulfur and silicon emission lines in SNR N132D are while that of the Fe He emission line is (Xrism2024N132D)
SNR N132D is the fifteenth CCSNR with an identified point-symmetric morphology attributed to the JJEM (Section 2). Presently, the point-symmetric morphology is the only property of CCSNe and CCSNRs that can clearly distinguish between the JJEM and the delayed neutrino explosion mechanism (Soker2024UnivReview). Other properties, such as light curves and nucleosynthesis, are similar but not identical between the two explosion mechanisms. However, the differences in these properties are not at a level that allows for a clear observational determination of the explosion mechanism (Soker2024UnivReview). The point-symmetrical morphology due to point-symmetric explosion is a robust prediction of the JJEM (although it will not be observed in all CCSNRs) that has no explanation in the neutrino-driven mechanism (e.g., SokerShishkin2025Vela). Some supporters of the neutrino-driven explosion mechanism have noted that their simulations do not produce jets (I attributed it to the limited ability of existing CCSN hydrodynamical numerical codes to handle magnetic field reconnection; Soker2025Learning). As a result, they often do not engage with or cite literature related to the JJEM model. This approach, where theoretical expectations (but only by some researchers) are used to dismiss observational evidence, contrasts with the standard scientific methodology in which models are tested and refined based on empirical data. In contrast, the JJEM model argues that the observed point symmetric morphologies in CCSNRs provide significant evidence that favors jet-driven processes over neutrino-driven explosions as the primary explosion mechanism. I therefore encourage the CCSNe research community to consider evaluating the JJEM on equal ground with the delayed neutrino explosion mechanism.
Acknowledgements
I thank Ealeal Bear, Dima Shishkin, and Amit Kashi for their many discussions and comments, which were relevant to the topic of this study. I thank an anonymous referee for useful comments. I thank the Charles Wolfson Academic Chair at the Technion for the support.